How many liters each of 12.5% acid solution and 20.25% acid solution must be mixed to obtain 100 liters of 17.5% acid solution?
(Please show the work.)

Answer by stanbon(75887)

How many liters each of 12.5% acid solution and 20.25% acid solution must be mixed to obtain 100 liters of 17.5% acid solution?
-----------------------
Equation:
acid + acid = acid
0.125x + 0.2025(100-x) = 0.175*100
----
Multiply thru by 10000 to get rid of the decimals:
1250x + 2025(100-x) = 1750*100
-775x = -275*100
x = 35.48 liters (amt of 12.5% solution in the mixture)
100-35.48 = 64.52 liters (amt of 20.25% solution in the mixture)
----------------------------
Cheers,
Stan H.